Cupid is hovering in the air when he shoots an arrow giving it a horizontal velocity of 14 m/s. If the arrow covers a horizontal displacement of 17 m before landing on the ground, from what height was the arrow shot from Cupid’s bow?


1
Expert's answer
2021-01-13T11:35:48-0500

Let's first find the time that the arrow takes to reach the ground:


"x=v_0t,""t=\dfrac{x}{v_0}=\dfrac{17\ m}{14\ \dfrac{m}{s}}=1.21\ s."

Finally, we can find the height from which the arrow was shot:


"y=\dfrac{1}{2}gt^2,""y=\dfrac{1}{2}\cdot 9.8\ \dfrac{m}{s^2}\cdot(1.21\ s)^2=7.2\ m."

Answer:

"y=7.2\ m."